OpenStudy (anonymous):

Express 6615 under a radical sign in simplest form.

OpenStudy (kirbykirby):

\(\sqrt{6615}=\sqrt{5*9*147}=3\sqrt{5*147}=3\sqrt{5*7*21}=3\sqrt{5*7*7*3}\) =\(3*7\sqrt{5*3}=21\sqrt{15}\)

OpenStudy (kirbykirby):

:)

OpenStudy (kirbykirby):

the easiest thing is just to try dividing with small numbers first. 6615 ends in a 5, so it's divisible by 5... etc.

OpenStudy (anonymous):

ok thanks :)